OpenStudy (anonymous):

r^2(cosθrsinθ)=1/2r^2sin(2θ) how are these equivalent?

myininaya (myininaya):

\[\text{ recall } \sin(2\theta)=2 \sin(\theta) \cos(\theta)\] \[\text{ so } \frac{1}{2}\sin( 2 \theta) =\frac{1}{2} \cdot 2\sin(\theta) \cos(\theta)=\sin(\theta) \cos(\theta)\]

OpenStudy (anonymous):

ahh..that is what is getting me.. thank you! awesome!!!

myininaya (myininaya):

np

OpenStudy (anonymous):

its the "Double Angle Formula!!!" lol